# SUSTAINABLE Platform: Seamless Smart Farming Integration Towards Agronomy Automation

### v1

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
The global agricultural sector faces mounting challenges [1], including rising food demand [2], climate variability [3], and increasing concerns over food safety and security [4]. Traditional farming methods often struggle to ensure consistent crop quality while minimizing resource waste and contamination risks. In response, the integration of Agriculture 4.0 [5]—and especially the application of Industry 4.0 technologies in farming [6]—has emerged as a transformative approach to achieving smart, automated and sustainable agronomy. By leveraging advancements such as the Internet of Things (IoT) [7], artificial intelligence (AI) [8], robotics [9], and big data analytics [10], modern farming systems can enhance precision, efficiency, and traceability across the agricultural value chain. Different challenges persist in the use and integration of diverse technologies towards Agriculture 4.0 paradigm [11].
@@
The present work presents H2020 SUSTAINABLE project [12], and its agricultural process management platform that actively engages farming ecosystem stakeholders – including agronomists, farmers, and cooperatives – in executing data-informed actions across the entire cultivation cycle, from planting to harvest. Unlike conventional systems that merely aggregate raw data, the platform operates at the information level, transforming inputs from IoT sensors, drone and satellite imagery, smart lab analyzers, and meteorological sources into actionable insights for stakeholders. By integrating ready-to-use solutions not as isolated data streams but as interconnected decision-support tools, the platform enables farmers and agronomists to plan, monitor, and adjust agricultural practices in real time. This approach ensures that stakeholders remain central to the process, leveraging technology to enhance precision, sustainability, and compliance without replacing human expertise.
@@ II Related Work - Integration of Smart Agriculture Technologies
xFarm combines weather forecasts with visualization to help monitor field conditions and plan treatments [13]. CropX uses soil sensors and Machine Learning (ML) based recommendations for irrigation and fertilization [14]. Climate FieldView focuses on combining satellite and ground-based data to track vegetation variability across the season [15]. OneSoil provides real-time NDVI analysis and zoning for input optimization [16].
Academic approaches complement these platforms. Kulatunga et al. use NDVI and topographic data for ML-driven zoning in vineyards [17], while Mazzia et al. enhance NDVI accuracy using UAV imagery and deep learning models tailored to viticulture [18]. Kumar et al. [19] review smart agriculture frameworks with emphasis on IoT-enabled data integration for sustainable practices.
@@ III-A Platform Conceptual Architecture
The Agricultural Processes Designer is a tool that simplifies farming process design targeting farm managers and agronomists. It provides users with a simple and intuitive GUI that helps them design agriculture-related processes or transform existing ones. Process design is based on a formal description language, the Business Process Model and Notation (BPMN) [20], a design based language adapted for the definition of business processes. The process includes sequence of the actions to be taken, actors to perform actions, and information to be collected and exchanged during task execution. The output of the modeler is a valid BPMN file in XML format ready to be executed in the platform’s execution engine, tailored to the needs of each farm.
@@ III-E Business Process Management Notation Execution Engine
This is the platform’s core component for orchestrating the agricultural enterprise’s operational plans through the design and execution of formal business processes. To model these workflows, we have adopted the Business Process Model and Notation 2.0 (BPMN 2.0) [20] standard. This specification was selected because it is the de facto industry solution, offering both straightforward translation into software components and complete independence from any particular technology stack.
Execution of these process models is managed by Camunda [21], a robust open-source engine for workflow and decision automation. While the conceptual design of a process can be performed with any BPMN 2.0-compliant software, the development of a deployable, executable artifact requires the use of the specialized Camunda Modeler. This tool is essential for linking the business model to the underlying technical implementation and deploying it to the engine.
@@ IV Platform Use Case
The SUSTAINABLE platform was evaluated by executing a use case scenario in a real-world farming environment. The scenario was conducted at a small, independent Greek winery that cultivates several native grape varieties [22] and involved modeling its daily and annual operational processes. These agricultural processes are supported by several external data systems. The Macro Expert Kft platform supplied weather, soil, and pest infection predictions, while Graniot [23] provided processed satellite imagery. When satellite data was insufficient, GAIA Robotics [24] gathered supplemental data using drones. Finally, DNA Phone [25] performed on-demand smart quality control (QC) with its portable devices. Following a thorough analysis, the vineyard’s complete operational workflow was documented in a BPMN diagram (Figure 2). Additionally, the necessary API callers were developed to enable communication with the third-party applications.

### v2

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. The definitive version was published in IEEE International Smart Cities Conference (ISC2) Prooceedings, DOI: 10.1109/ISC266238.2025.11293287.
@@ I Introduction
The global agricultural sector faces mounting challenges [vos2019global], including rising food demand [falcon2022rethinking], climate variability [bhatti2024global], and increasing concerns over food safety and security [barrett2021overcoming]. Traditional farming methods often struggle to ensure consistent crop quality while minimizing resource waste and contamination risks. In response, the integration of Agriculture 4.0 [rose2018agriculture]—and especially the application of Industry 4.0 technologies in farming [batra2024industrial]—has emerged as a transformative approach to achieving smart, automated and sustainable agronomy. By leveraging advancements such as the Internet of Things (IoT) [ojha2021internet], artificial intelligence (AI) [ahmad2024ai], robotics [yepez2023mobile], and big data analytics [chergui2022data], modern farming systems can enhance precision, efficiency, and traceability across the agricultural value chain. Different challenges persist in the use and integration of diverse technologies towards Agriculture 4.0 paradigm [alexopoulos2023complementary].
@@
The present work presents H2020 SUSTAINABLE project [sustainable], and its agricultural process management platform that actively engages farming ecosystem stakeholders – including agronomists, farmers, and cooperatives – in executing data-informed actions across the entire cultivation cycle, from planting to harvest. Unlike conventional systems that merely aggregate raw data, the platform operates at the information level, transforming inputs from IoT sensors, drone and satellite imagery, smart lab analyzers, and meteorological sources into actionable insights for stakeholders. By integrating ready-to-use solutions not as isolated data streams but as interconnected decision-support tools, the platform enables farmers and agronomists to plan, monitor, and adjust agricultural practices in real time. This approach ensures that stakeholders remain central to the process, leveraging technology to enhance precision, sustainability, and compliance without replacing human expertise.
@@ II Related Work - Integration of Smart Agriculture Technologies
xFarm combines weather forecasts with visualization to help monitor field conditions and plan treatments [xfarm]. CropX uses soil sensors and Machine Learning (ML) based recommendations for irrigation and fertilization [cropx]. Climate FieldView focuses on combining satellite and ground-based data to track vegetation variability across the season [fieldview]. OneSoil provides real-time NDVI analysis and zoning for input optimization [onesoil].
Academic approaches complement these platforms. Kulatunga et al. use NDVI and topographic data for ML-driven zoning in vineyards [kulatunga2024], while Mazzia et al. enhance NDVI accuracy using UAV imagery and deep learning models tailored to viticulture [mazzia2020]. Kumar et al. [kumar2024] review smart agriculture frameworks with emphasis on IoT-enabled data integration for sustainable practices.
@@ III-A Platform Conceptual Architecture
The Agricultural Processes Designer is a tool that simplifies farming process design targeting farm managers and agronomists. It provides users with a simple and intuitive GUI that helps them design agriculture-related processes or transform existing ones. Process design is based on a formal description language, the Business Process Model and Notation (BPMN) [bpmnspec:online], a design based language adapted for the definition of business processes. The process includes sequence of the actions to be taken, actors to perform actions, and information to be collected and exchanged during task execution. The output of the modeler is a valid BPMN file in XML format ready to be executed in the platform’s execution engine, tailored to the needs of each farm.
@@ III-E Business Process Management Notation Execution Engine
This is the platform’s core component for orchestrating the agricultural enterprise’s operational plans through the design and execution of formal business processes. To model these workflows, we have adopted the Business Process Model and Notation 2.0 (BPMN 2.0) [bpmnspec:online] standard. This specification was selected because it is the de facto industry solution, offering both straightforward translation into software components and complete independence from any particular technology stack.
Execution of these process models is managed by Camunda [camunda:online], a robust open-source engine for workflow and decision automation. While the conceptual design of a process can be performed with any BPMN 2.0-compliant software, the development of a deployable, executable artifact requires the use of the specialized Camunda Modeler. This tool is essential for linking the business model to the underlying technical implementation and deploying it to the engine.
@@ IV Platform Use Case
The SUSTAINABLE platform was evaluated by executing a use case scenario in a real-world farming environment. The scenario was conducted at a small, independent Greek winery that cultivates several native grape varieties [orfanos:online] and involved modeling its daily and annual operational processes. These agricultural processes are supported by several external data systems. The Macro Expert Kft platform supplied weather, soil, and pest infection predictions, while Graniot [graniot:online] provided processed satellite imagery. When satellite data was insufficient, GAIA Robotics [gaiarobotics:online] gathered supplemental data using drones. Finally, DNA Phone [dnaphone:online] performed on-demand smart quality control (QC) with its portable devices. Following a thorough analysis, the vineyard’s complete operational workflow was documented in a BPMN diagram (Figure 2). Additionally, the necessary API callers were developed to enable communication with the third-party applications.